OpenStudy (anonymous):

solve for T T(45)=(75)+(110)e^(-0.027*45)

OpenStudy (mertsj):

-.027(45)=-1.215 e^-1.215=.29671 110(e^-1.215)=110(.29671)=32.63810 75+110e^-1.215=75+32.63810=107.637